find the function of x

f(x) = x^2 + 3x / 2x^2 + 6x + 4
1
Expert's answer
2011-02-09T19:27:29-0500
It's not clear what to do with this function, we can factorise it and find zeroes of it:
f(x) = (x^2 + 3x )/( 2x^2 + 6x + 4) = x(x+3)/2(x+1)(x+2)
f(x) = 0 at x=0 and x=3, f(x) doesn't exist at x = -1 and x = -2.